OpenStudy (anonymous):

f(x)=2^3x+5 finding f^-1 (x) ?

OpenStudy (watchmath):

Is that \(f(x)=2^{3x}+5\)?

OpenStudy (anonymous):

correcto

OpenStudy (watchmath):

\(y=2^{3x}+5\) \(y-5=2^{3x}\) \(\log_2(y-5)=3x\) \(x=\frac{1}{3}\log_2(y-5)\) So \(f^{-1}(x)=\frac{1}{3}\log_2(x-5)\)

OpenStudy (anonymous):

wow! thanx